# SCILLA: SurfaCe Implicit Learning for Large Urban Area, a volumetric hybrid solution

### v2

@@ Abstract
Neural implicit surface representation methods have recently shown impressive 3D reconstruction results. However, existing solutions struggle to reconstruct urban outdoor scenes due to their large, unbounded, and highly detailed nature. Hence, to achieve accurate reconstructions, additional supervision data such as LiDAR, strong geometric priors, and long training times are required. To tackle such issues, we present SCILLA, a new hybrid implicit surface learning method to reconstruct large driving scenes from 2D images. SCILLA’s hybrid architecture models two separate implicit fields: one for the volumetric density and another for the signed distance to the surface. To accurately represent urban outdoor scenarios, we introduce a novel volume-rendering strategy that relies on self-supervised probabilistic density estimation to sample points near the surface and transition progressively from volumetric to surface representation. Our solution permits a proper and fast initialization of the signed distance field without relying on any geometric prior on the scene, compared to concurrent methods. By conducting extensive experiments on four outdoor driving datasets, we show that SCILLA can learn an accurate and detailed 3D surface scene representation in various urban scenarios while being two times faster to train compared to previous state-of-the-art solutions.
Keywords:
Neural Implicit Surface Reconstruction Radiance Fields Mesh Reconstruction
@@
Figure 1: We introduce SCILLA, a novel hybrid implicit learning method capable to accurately reconstruct large driving scenes from 2D input images solely. Extensive experiments on four popular driving datasets show the superiority of SCILLA’s mesh (left) over previous state-of-the-art methods such as StreetSurf[11] (right). Figure shows both original and textured mesh obtained from both methods.
@@ 1 Introduction
Achieving accurate 3D reconstruction from multi-view-consistent images has been an important topic since decades in computer vision and computer graphics communities. A 3D representation of the perceived environment is fundamental for addressing many tasks, such as scene re-lighting [41], scene editing [42] and 3D object insertion for artificial data augmentation [36].
@@
Recently, neural implicit surface representation methods[34, 39] have shown impressive 3D surface reconstruction performance from monocular images, achieving highly detailed and precise mesh representations. Such methods are originally inspired by Neural Radiance Fields (NeRF) [21], and reframed for multi-view surface reconstruction. NeRF introduced volume rendering for Novel View Synthesis (NVS) by sampling points along camera rays. These points, along with their viewing directions, are fed into a Multi-Layer Perceptron (MLP) to predict their respective colors and densities. The final image is rendered by blending these predicted values using alpha compositing. Conversely, neural surface representation methods aim to approximate the surface using a learnable Signed Distance Function (SDF) [34, 39]. Unlike NeRF, the outputs of the MLP are colors and SDF values instead of colors and volumetric densities, which are later blended using a modified alpha-compositing equation adapted to SDF. An accurately reconstructed surface is represented as a zero-level set of the SDF. Assuming the objects lying at the center of the scene, the SDF is first initialized to form a spherical shape to bound the region of interest. Although current methods achieve impressive reconstruction results, they are mostly applied to simple, bounded and object-centric scenes. Hence, applying such methods to unbounded urban scenarios results in sub-optimal reconstruction given their significantly different nature. To overcome such limitations, implicit surface reconstruction approaches commonly rely on LiDAR supervision [25, 36] or strong geometric priors [11] and require longer training times to produce qualitative outdoor scene reconstruction meshes from 2D images.
@@
To address those challenges, we propose SCILLA, a hybrid approach to accurately initialize a signed distance field of an unbounded area and represent complex surfaces from urban outdoor scenarios with high accuracy and geometric precision while requiring half of the training time compared to other state-of-the-art solutions such as StreetSurf [11]. We propose a hybrid architecture that models two separate implicit fields: one representing the volumetric density of the scene and another one representing the signed distance to the surface. SCILLA’s novel volume rendering approach relies on a self-supervised density estimator which permits to sample points near the surface along with a progressive sampling strategy that smoothly transitions from a volumetric to a surface representation, drastically reducing the model convergence time. Inspired by previous state-of-the-art works [11, 40], we supervise surface normals by relying on normal cues predicted from a pre-trained monocular estimator [8]. We extensively validate our solution on KITTI-360 [18], Pandaset [38], Waymo Open Dataset [29], and nuScenes [3]. By comparing the quality of the extracted meshes from the signed distance field, we show that our method can efficiently reconstruct urban outdoor scenarios showing SoTA accuracy. To summarize, the main contributions of our work are:
@@
  • A novel implicit hybrid learning method for surface reconstruction on large outdoor driving scenes,
@@ 2 Related work
NeRF-based methods can achieve impressive NVS results but fail to accurately reconstruct scene surfaces given the lack of constraints or priors on the surface when modelling the scene as a density volume. Neural surface reconstruction approaches on the other hand, explicitly represent the surface as an SDF and propose various adaptations to train such methods through differentiable volume rendering [34, 39]. NeuS [34] proposes a novel SDF-based rendering equation that relies on SDF values to compute the alpha blending factor used in volume rendering instead of densities as presented in the original NeRF [21]. Nevertheless, NeuS’s original contribution is designed mainly for object-centric datasets and takes long training times. Some recent works aim to reduce training times [35] or reconstruct large-scale highly detailed landmark scenes [17]. However, such approaches are not tailored for large unbounded driving scenarios and still require relatively long training times. In contrast, our method, is built upon NeuS’s formulation and it is designed to reliably represent urban outdoor scenes.
@@
Urban outdoor scene reconstruction is one of the main subjects contributing to resolving many downstream applications, such as inverse rendering and object insertion [36]. Latest state-of-the-art approaches rely on SDF-based surface reconstruction for urban outdoor scenes, to disentangle the scene’s geometry from other scene parameters.
Existing SoTA methods relying on neural implicit surface reconstruction are mainly guided by LiDAR [36, 25], limiting their large-scale deployment.
One of the few existing approaches that tackle the subject of surface reconstruction of urban outdoor scenes is StreetSurf[11]. The authors of StreetSurf present a solution that disentangles close-range view from distant-range view, alongside a signed distance field initialization technique to match the road surface according to the ego poses of the vehicle. This surface initialization approach is rough, not straightforward, and heavily depends on the ego poses and the assumption that the road and the sidewalk are at the same height in every part of the scene. Additionally, StreetSurf relies on the assumption that self-driving scenes are mostly long and narrow and showcase reconstruction of the scene from only one public dataset (i.e. Waymo Open Dataset [29]). In reality, driving datasets frequently contain scenarios where the scenes are short and curvy, hence, disentangling the close and far range in those situations becomes more challenging. Instead, our solution is more adaptive in comparison to StreetSurf as it does not rely on strong assumptions about the scene’s topology. Furthermore, SCILLA is also more compact and faster to train as the entire scene is represented by a single hybrid implicit field.
@@
Some state-of-the-art methods have already investigated the use of hybrid rendering solutions relying on both volumetric and surface composing. Wang et al. [37] introduced a hybrid rendering method based on NeuS [34] that uses a spatially varying kernel, which changes according to the fuzziness of the surface. Based on the kernel, a narrow mesh envelope is extracted around the surface, which then enables ray casting only inside this envelope. This solution helps accelerate both rendering time and quality. Other methods [12, 13, 30] focus on a hybrid representation that both relies on an explicit representation (meshes) and an SDF representation. The hybrid volume-mesh representation enables efficient rendering and less resourceful storage while outputting an implicit geometrical representation. In addition to fast training times, this representation also benefits from easy integration into computer graphics engines. NeuSG [5] combines 3D Gaussian Splatting [15] with NeuS to obtain an accurate Gaussian cloud close to the surface that serves as a direct supervision signal during NeuS training. The aforementioned solutions mainly focus on improving the rendering speed and are better fitted for small scenes with a hybrid representation. In our proposal, we jointly train a hybrid model to benefit from the fast training of a density-based approach along with an accurate geometric representation obtained thanks to the signed distance field learning.
@@ 3.1 Preliminaries
3.1.1 NeRF’s Implicit Volumetric Representation.
NeRF [21] takes as input the 3D positions $x_{i}\in\mathbb{R}^{3}$ and the viewing direction $d\in\mathbb{R}^{2}$ of sampled points along each camera ray $r$ and predicts their corresponding densities $\sigma_{i}\in\mathbb{R}$ and colors $c_{i}\in\mathbb{R}^{3}$. This representation of the scene is approximated using a Multi-Layer Perceptron (MLP). Finally, volume rendering is used to alpha-composite the colors of the samples along each ray and render the final pixel color $\hat{C}(r)\in\mathbb{R}^{3}$:
@@
3.1.2 SDF representation: NeuS.
For an accurate surface-based representation of a 3D scene, NeuS [34] presented a different formulation of the NeRF’s volume rendering equation. They represent the scene as a signed distance function, where the surface is the zero-level set of such function: $S=\left\{x\in\mathbb{R}^{3}|f(x)=0\right\}$, where $f(x)$ is the signed distance value at a given position $x$. To efficiently represent surfaces, NeuS adopted a volume rendering formulation introducing a density distribution function which is formulated as the derivative of the sigmoid function $\Phi_{s}(x)=(1+e^{-sx})^{-1}$, that is: $\phi_{s}(x)=se^{-sx}/(1+e^{-sx})^{2}$. The standard deviation of $\phi_{s}$ is $1/s$ and, as the network converges to an accurate surface representation, this parameter moves toward zero. Following NeuS’s formulation, the volumetric alpha used in Eq. 1 is replaced by the following equation to approximate the color $\hat{C}(r)$:
@@ 3.2 SCILLA’s hybrid architecture
Unlike SoTA methods that singularly represent either a density or an SDF field, our hybrid method SCILLA learns the two different fields simultaneously, as illustrated in Fig 3. SCILLA expresses the scene using two functions, $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$, encoded respectively by two MLPs. The input 3D position $x$ and viewing direction $d$ are embedded respectively via a positional and directional embedding into $\text{emb}_{x}(x)$ and $\text{emb}_{d}(d)$.
@@
[Uncaptioned image] Figure 3: SCILLA architecture – Our method uses two MLP functions $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$ to output SDF values and density ($\sigma$ and $f(x)$, respectively) along with color values given an input sample. We inspire from NeRF and design $\mathcal{F}_{\Theta}^{c}$ to output the color given the lattent vector $h$ outputted by $\mathcal{F}_{\Theta}^{h}$, the viewing direction ($d$) and the normal vector $\overrightarrow{n}$ obtained from the gradient of the SDF
[Uncaptioned image] Figure 4: Sample attribution during hybrid stage – Left: we start by assigning samples with the lowest density, i.e farthest away from the real surface, to alpha values computed from the SDF (green dots) and the rest to alpha values computed from estimated density (red dots). Middle, right: as the training advances, the proportion of samples associated with SDF values increases and the signed distance field (dotted green line) can be estimated more precisely.
@@ 3.3 SCILLA’s volume rendering
3.3.1 Probabilistic density estimation.
We rely on probabilistic density estimation along the ray for a more targeted sampling strategy that samples fewer points and accelerates the overall method’s training. We sample $N$ points guided by the estimated density along the ray, where higher density indicates the presence of the surface. Given these samples, we compute using SCILLA’s architecture the alpha values and perform volume rendering to obtain the final color associated with the ray, which is compared to the ground truth value to train our model:
@@
The strategy of associating a sample to either a volumetric or surface-based representation is discussed in the next section. The weights $w_{i}$ associated with each sample (Eq. 1) serve as a supervision signal to our probabilistic density estimator. Hence, we can refine in a self-supervised manner the prediction of the expected density along the ray. We rely on Mip-NeRF 360’s [1] proposal networks as our volumetric density estimator and use their introduced proposal loss, $\mathcal{L}_{\text{prop}}$, to refine density estimation from the weights $w_{i}$ of our main model.
3.3.2 Volumetric to surface representation.
Starting from a pure volumetric density training ($\alpha_{i}=\alpha_{i}^{v}$), we introduce a progressive transition from a volumetric to an SDF representation. Our approach is based on the premise that estimating the density $\sigma_{i}$ at a specific location requires less complex geometric understanding in contrast to estimating the signed distance to the closest point to the surface (i.e. $f(x_{i})$). By initially leveraging a pure volumetric density representation during training, we shorten the convergence time of our density estimator. Our progressive transitioning strategy can be outlined as follows:
@@
3.3.3 Hybrid stage regularization.
Composing alpha values obtained from our two different geometric representations is not straightforward as the learning dynamic of the density and the SDF are not the same. As mentioned earlier, learning a volumetric representation of the scene based only on spatial density is easier and faster to learn compared to learning an arbitrary signed distance field of a complex outdoor scene. To facilitate the convergence of our representation, we introduce two regularization methods during the hybrid stage: a probability-based sampling attribution and an SDF-gradient normalization.
Probability-based samples attribution. To assign the alpha value based on density or SDF to a sample, we initially start with a random selection for each ray. We noticed that this association strategy is not suitable: the signed distance field could not be properly initialized during the hybrid stage. To solve this problem, we introduce a probability-based sample association. We assign $\alpha_{i}^{f}$ to samples with the lowest predicted density to facilitate the learning of the signed distance field. Following this sample attribution strategy, we can estimate a coarse surface during the first iterations and refine the fine geometrical details as the training progresses, as illustrated in the toy example in Fig. 4.
SDF-gradient normalization. Volumetric representations based on density estimation tend to quickly approximate saturated alpha values ($\alpha_{i}^{v}$ being either 00 or $1$), while alpha values computed from the SDF converge slower. Because $\alpha_{i}^{v}$ and $\alpha_{i}^{f}$ are composed jointly during the hybrid stage, we noticed that surface representation compensates for this gap by predicting large SDF gradient values so that $\alpha_{i}^{f}$ aligns with $\alpha_{i}^{v}$. Indeed, from Eq. 3 a solution to saturate alpha towards either 1 or 0, is to predict $f(p_{i+1})\ll f(p_{i})$ or $f(p_{i+1})\gg f(p_{i})$, respectively. During the forward pass, $f(p_{i})$ and $f(p_{i+1})$ are not directly predicted by the model but rather computed using:
with $\theta$ being the angle between the direction of $\nabla f(x_{i})$ and the ray direction $d$. Practically, we do not formally compute $\theta$ but we rather approximate its cosine with $\cos({\theta})=\nabla f(x_{i})\cdot d$, assuming both vectors are unit norm. By predicting $\left\|\nabla f(x_{i})\right\|_{2}\gg 1$, $\alpha_{i}^{f}$ can be easily saturated and follow the distribution of $\alpha_{i}^{v}$ without learning a proper signed distance field. To address this, we simply normalize the SDF gradient before the cosine computation to prevent the gradient from compensating the alpha distribution difference. It is important to notice that even if the eikonal loss used for training (see Sec. 3.4.3) is supposed to encourage the network to model a signed distance function with spatial derivative of unitary norm, we found that numerically normalizing the gradient during our hybrid stage is essential.

### v3

@@ Abstract
Neural implicit surface representation methods have recently shown impressive 3D reconstruction results. However, existing solutions struggle to reconstruct driving scenes due to their large, unbounded, and highly complex nature. Hence, to achieve accurate reconstructions, additional supervision data such as LiDAR, strong geometric priors, and long training times are required. To tackle such limitations, we present SCILLA, a new hybrid implicit surface learning method to reconstruct large driving scenes from 2D images. SCILLA’s hybrid architecture models two separate implicit fields: one representing the volumetric density of the scene, and another one representing the signed distance to the surface. To accurately reconstruct urban outdoor driving scenarios, we introduce a novel volume-rendering strategy that relies on self-supervised probabilistic density estimation to sample points near the surface and transition progressively from volumetric to surface representation. Our solution permits a proper and fast initialization of the signed distance field without relying on any geometric prior on the scene, compared to concurrent methods. By conducting extensive experiments on four outdoor driving datasets, we show that SCILLA can learn an accurate and detailed 3D surface scene representation in various driving scenarios while being two times faster to train compared to previous state-of-the-art solutions.
@@
Figure 1: We introduce SCILLA, a novel hybrid implicit learning method capable to accurately reconstruct large driving scenes from 2D RGB input images solely. Extensive experiments on four popular driving datasets show the superiority of SCILLA’s mesh (left) over previous state-of-the-art methods such as StreetSurf [11] (right). Figure shows both original and textured mesh obtained from both methods.
@@ 1 Introduction
Achieving accurate 3D reconstruction from multi-view-consistent images has been an important topic in computer vision and computer graphics communities for decades. A 3D representation of the perceived environment is fundamental for many tasks, such as scene re-lighting [44], scene editing [45] and 3D object insertion for artificial data augmentation [39].
@@
Recently, neural implicit surface representation methods [37, 42] have shown impressive 3D surface reconstruction performance from monocular images, achieving highly detailed and precise mesh representations. Such methods are originally inspired by Neural Radiance Fields (NeRF) [23], and re-framed for multi-view surface reconstruction. NeRF introduced volume rendering for Novel View Synthesis (NVS) by sampling points along camera rays. These points, along with their viewing directions, are fed into a Multi-Layer Perceptron (MLP) to predict their respective colors and densities. The final image is rendered by blending these predicted values using alpha compositing. Conversely, neural surface representation methods aim to approximate the surface using a learnable Signed Distance Function (SDF) [37, 42]. Unlike NeRF, the outputs of the MLP are colors and SDF values instead of colors and volumetric densities, which are later blended using a modified alpha-compositing equation adapted to SDF. An accurately reconstructed surface is represented as a zero-level set of the SDF. Assuming the objects lying at the center of the scene, the SDF is first initialized to form a spherical shape to bound the region of interest. Although current methods achieve impressive reconstruction results, they are mostly applied to simple, bounded and object-centric or landmarks scenes. Applying such methods to unbounded driving scenarios results in sub-optimal reconstruction given their significantly different nature. To overcome such limitations, implicit surface reconstruction of large outdoor driving scenes commonly relies on LiDAR supervision [27, 39] or strong geometric priors [11]. These approaches require longer training time to produce qualitative outdoor scene reconstruction meshes from 2D images.
@@
To address those challenges, we propose SCILLA, a hybrid approach to accurately initialize a signed distance field of an unbounded area and represent complex surfaces from urban outdoor driving scenarios with high accuracy and geometric precision, while requiring half of the training time compared to other state-of-the-art solutions such as StreetSurf [11]. We propose a hybrid architecture that models two separate implicit fields: one representing the volumetric density of the scene and another one representing the signed distance to the surface. SCILLA’s novel volume rendering approach relies on a self-supervised density estimator which permits to sample points near the surface along with a progressive sampling strategy that smoothly initializes and transitions from a volumetric to a surface representation, drastically reducing the model convergence time. Inspired by previous state-of-the-art works [11, 43], we supervise surface normals by relying on normal cues predicted from a pre-trained monocular estimator [8]. We extensively validate our solution on KITTI-360 [20], Pandaset [41], Waymo Open Dataset [32], and nuScenes [3]. By comparing the quality of the extracted meshes from the signed distance field, we show that our method can efficiently reconstruct urban outdoor scenarios showing SoTA accuracy. To summarize, the main contributions of our work are:
@@
  • A novel implicit hybrid learning method to progressively initialize the signed distance field for surface reconstruction on large outdoor driving scenes,
@@ 2 Related work
NeRF-based methods can achieve impressive NVS results but fail to accurately reconstruct scene surfaces given the lack of constraints or priors on the surface when modelling the scene as a density volume. Neural surface reconstruction approaches on the other hand, explicitly represent the surface as an SDF and propose various adaptations to train such methods through differentiable volume rendering [37, 42]. NeuS [37] proposes a novel SDF-based rendering equation that relies on SDF values to compute the alpha blending factor used in volume rendering instead of densities as presented in the original NeRF [23]. Nevertheless, NeuS’s original contribution is designed mainly for object-centric datasets and requires a long training time. Some recent works aim to reduce the training time [38] or reconstruct large-scale highly detailed landmark scenes [19]. However, such approaches are not tailored for large unbounded driving scenarios and still require hours of training. In contrast, our method is built upon NeuS’s formulation and it is designed to reliably represent urban outdoor driving scenes.
@@
Urban outdoor scene reconstruction is one of the main subjects contributing to resolving many downstream applications, such as inverse rendering and object insertion [39]. Latest SoTA approaches rely on SDF-based surface reconstruction for urban outdoor scenes, to disentangle the scene’s geometry from other scene parameters. Neuralangelo [19], a SoTA method for outdoor landmarks reconstruction, requires many overlapping images and the region of interest of the scene to be bounded to initialize the SDF with a spherical shape. Few methods consider acquisition from a moving camera rig for driving sequences surface reconstruction. Existing solutions are mainly guided by LiDAR such as FEGR [39] and Urban Radianc Fields [27], limiting their large-scale deployment. StreetSurf [11], the most related method to our proposal, presents a solution that disentangles close-range view from distant-range view, alongside a signed distance field initialization technique to match the road surface according to the ego poses of the vehicle. This surface initialization approach is rough, not straightforward, and heavily depends on the ego poses and the assumption that the road and the sidewalk are at the same height in every part of the scene. Additionally, StreetSurf relies on the assumption that self-driving scenes are mostly long and narrow and showcase reconstruction of the scene from only one public dataset (i.e. Waymo Open Dataset [32]). In reality, driving datasets frequently contain scenarios where the scenes are short and curvy, hence, disentangling the close and far range in those situations becomes more challenging. Instead, our solution is more adaptive in comparison to StreetSurf as it does not rely on strong assumptions about the scene’s topology. Furthermore, SCILLA is also more compact and faster to train as the entire scene is represented by a single hybrid implicit field.
@@
Some SoTA methods have already investigated the use of hybrid rendering solutions relying on both volumetric and surface composing. Wang et al. [40] introduced a hybrid rendering method based on NeuS [37] that uses a spatially varying kernel, which changes according to the fuzziness of the surface. Based on the kernel, a narrow mesh envelope is extracted around the surface, which then enables ray casting only inside this envelope. This solution helps accelerate both rendering time and quality. Other methods [12, 14, 33] focus on a hybrid representation that both relies on an explicit representation (meshes) and an SDF representation. The hybrid volume-mesh representation enables efficient rendering and less resourceful storage while outputting an implicit geometrical representation. In addition to fast training times, this representation also benefits from easy integration into computer graphics engines. NeuSG [5] combines 3D Gaussian Splatting [17] with NeuS to obtain an accurate Gaussian cloud close to the surface that serves as a direct supervision signal during NeuS training. The aforementioned solutions mainly focus on improving the rendering speed and are better fitted for small scenes with a hybrid representation. In our proposal, we jointly train a hybrid model to benefit from the fast training of a density-based approach along with an accurate geometric representation obtained thanks to the signed distance field learning.
@@ 3.1 Preliminaries
NeRF’s Implicit Volumetric Representation. NeRF [23] takes as input the 3D positions $x_{i}\in\mathbb{R}^{3}$ and the viewing direction $d\in\mathbb{R}^{2}$ of sampled points along each camera ray $r$ and predicts their corresponding densities $\sigma_{i}\in\mathbb{R}$ and colors $c_{i}\in\mathbb{R}^{3}$. This representation of the scene is approximated using a Multi-Layer Perceptron (MLP). Finally, volume rendering is used to alpha-composite the colors of the samples along each ray and render the final pixel color $\hat{C}(r)\in\mathbb{R}^{3}$:
@@
SDF representation with NeuS. For an accurate surface-based representation of a 3D scene, NeuS [37] presented a different formulation of the NeRF’s volume rendering equation. They represent the scene as a signed distance function, where the surface is the zero-level set of such function: $S=\left\{x\in\mathbb{R}^{3}|f(x)=0\right\}$, where $f(x)$ is the signed distance value at a given position $x$. NeuS adoptes a volume rendering formulation introducing a density distribution function which is formulated as the derivative of the sigmoid function $\Phi_{s}(x)=(1+e^{-sx})^{-1}$, that is: $\phi_{s}(x)=se^{-sx}/(1+e^{-sx})^{2}$. The standard deviation of $\phi_{s}$ is $1/s$ and, as the network converges to an accurate surface representation, this parameter moves toward zero. Following NeuS’s formulation, the volumetric alpha used in Eq. 1 is replaced by the following equation to approximate the color $\hat{C}(r)$:
@@ 3.2 SCILLA’s hybrid architecture
Unlike SoTA methods that singularly represent either a density or an SDF field, our hybrid method SCILLA learns the two different fields simultaneously, as illustrated in Fig. 3. SCILLA expresses the scene using two functions, $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$, encoded respectively by two MLPs. The input 3D position $x$ and viewing direction $d$ are embedded respectively via a positional and directional embedding into $\text{emb}_{x}(x)$ and $\text{emb}_{d}(d)$.
@@
Figure 3: SCILLA architecture – Our method uses two MLP functions $\mathcal{F}_{\Theta}^{h}$ and $\mathcal{F}_{\Theta}^{c}$ to output SDF values and density ($\sigma$ and $f(x)$, respectively) along with color values given an input sample. We inspire from NeRF and design $\mathcal{F}_{\Theta}^{c}$ to output the color given the lattent vector $h$ outputted by $\mathcal{F}_{\Theta}^{h}$, the viewing direction ($d$) and the normal vector $\overrightarrow{n}$ obtained from the gradient of the SDF.
@@ 3.3 SCILLA’s volume rendering
Probabilistic density estimation. Probabilistic density estimation along the ray is employed for a more targeted sampling strategy that samples fewer points and accelerates the overall method’s training time. Guided by the estimated density, $N$ points are sampled along the ray, where higher density indicates the presence of the surface. Using SCILLA’s architecture, alpha values are computed for these samples, and volume rendering is performed to obtain the final color associated with the ray. This resulting color is then compared to the ground truth value to train our model.
@@
The strategy of associating a sample to either a volumetric or surface-based representation is discussed in the next section.
The weights $w_{i}$ associated with each sample (Eq. 1) serve as a supervision signal to our probabilistic density estimator. Hence, we can refine in a self-supervised manner the prediction of the expected density along the ray. We rely on Mip-NeRF 360’s [1] proposal networks as our volumetric density estimator and use their introduced proposal loss, $\mathcal{L}_{\text{prop}}$, to refine density estimation from the weights $w_{i}$ of our main model.
SDF progressive initialization: Volumetric to surface representation. Starting from a pure volumetric density training ($\alpha_{i}=\alpha_{i}^{v}$), we introduce a progressive initialization of the signed distance field, that transitions from a volumetric to an SDF representation. Our approach is based on the premise that estimating the density $\sigma_{i}$ at a specific location requires less complex geometric understanding in contrast to estimating the signed distance to the closest point to the surface (i.e. $f(x_{i})$). 3D geometry represented as density field contains less information than SDF field: the former contains occupancy information (which can be considered as binary) whereas the latter represents in addition continuous distances to the surface. Therefore, learning a density field is simpler compared to the SDF field (see Fig. 6 of our ablation section). By initially leveraging a pure volumetric density representation during training, we shorten the convergence time of our density estimator.
Our progressive initialization strategy can be outlined as follows:
@@
Hybrid stage regularization. Composing alpha values obtained from our two different geometric representations is not straightforward as the learning dynamic of the density and the SDF are not the same. As mentioned earlier, learning a volumetric representation of the scene based only on spatial density is easier and faster to learn compared to learning an arbitrary signed distance field of a complex outdoor scene. To facilitate the convergence of our representation, we introduce two regularization methods: probability-based sampling attribution and SDF-gradient normalization.
Probability-based samples attribution. To assign the alpha value based on density or SDF to a sample, we initially start with a random selection for each ray. We noticed that this association strategy is not suitable: the signed distance field could not be properly initialized during the hybrid stage. To solve this problem, we introduce a probability-based sample association. We assign $\alpha_{i}^{f}$ to samples with the highest predicted density to initialize the signed distance field. It facilitates the learning of the signed distance field by limiting the range of predicted values, as all the points are sampled close to a surface.
SDF-gradient normalization. During the hybrid stage, where $\alpha_{i}^{v}$ and $\alpha_{i}^{f}$ are jointly composed, we observed that the surface representation tends to predict large SDF gradient values to align $\alpha_{i}^{f}$ with $\alpha_{i}^{v}$ (we refer to our supplementary materials for further details). To address this, we normalize the SDF gradient in the $\alpha_{i}^{f}$ computation to prevent it from compensating for differences in the alpha distribution.